# Sylvester-Preconditioned Adaptive-Rank Implicit Time Integrators for Advection-Diffusion Equations with Variable Coefficients

### v2

@@ 3.4 Truncation, Efficient Residual Computation
Following [14], we implement a cost-effective way to compute the residual using the identity:
@@
$={\left\|R_{U}\begin{bmatrix}-\tilde{B}_{1}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&{S_{1,\epsilon}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&-\Delta t\,(I_{\mathcal{R}_{\text{ranks}}}\otimes{S_{1,\epsilon}})\end{bmatrix}R_{V}^{\top}\right\|}\leq\epsilon_{\texttt{tol}},$
@@
{_$\{\_\ ,R_{U}\}$ $=\texttt{QR}\left(\begin{bmatrix}{U}_{1}&{U}_{1,\epsilon}&{T_{1}}{U}_{1,\epsilon}&\Phi_{1}{U}_{1,\epsilon}&{T_{3}}{U}_{1,\epsilon}&\Sigma_{1}{U}_{1,\epsilon}\end{bmatrix}\right),$ (24)
{_$\{\_\ ,R_{V}\}$ $=\texttt{QR}\left(\begin{bmatrix}{V}_{1}&{V}_{1,\epsilon}&\Phi_{2}{V}_{1,\epsilon}&{T_{2}}{V}_{1,\epsilon}&\Sigma_{2}{V}_{1,\epsilon}&{T_{4}}{V}_{1,\epsilon}\end{bmatrix}\right).$
@@ 3.5 Generalization to Multi-rank Coefficients
[…]
@@
We summarize the pseudo-algorithm for multi-rank coefficients in Algorithm 4. The tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, $\epsilon_{\texttt{tol}}$ in the algorithm are set as follows. First, we choose the residual tolerance for xKrylov outer iterations, $\epsilon_{\texttt{tol}}$. To prevent rapid growth of the xKrylov subspaces, we choose the Krylov basis truncation threshold as $\epsilon_{\kappa}\leq\epsilon_{\texttt{tol}}$, typically two to three orders of magnitude smaller. The GMRES-ACS tolerance for the inner reduced system solve, $\epsilon_{\texttt{GMRES}}$, is initially set to be less than $\epsilon_{\texttt{tol}}$ for the first outer-basis iteration. It is subsequently adjusted to be two to three orders of magnitude smaller than the relative residual error from the previous xKrylov outer iteration. This adjustment ensures that errors from the inner solve do not pollute the accuracy of the outer solve. The SVD truncation threshold $\epsilon$ is chosen to be one to three orders of magnitude smaller than the tolerance $\epsilon_{\text{tol}}$. Since solution truncation occurs before residual evaluations, the truncation error in the solution should remain below the residual tolerance to prevent residual pollution by over-truncation.
@@ 7        Orthogonalize and truncate to tolerance $\epsilon_{\kappa}$;
17        Compute {_$\{\_,R_{U}\}=\texttt{QR}\left([U_{1},U_{1,\epsilon},T_{1,1:\ell_{x}}U_{1,\epsilon},\Phi_{1:\ell_{y}}^{1,y}U_{1,\epsilon},T_{3,1:k_{x}}U_{1,\epsilon},\Sigma_{1:k_{y}}^{1,y}U_{1,\epsilon}]\right)$ and
18        {_$\{\_,R_{V}\}=\texttt{QR}\left([V_{1},V_{1,\epsilon},\Phi_{1:\ell_{x}}^{2,x}V_{1,\epsilon},T_{2,1:\ell_{y}}V_{1,\epsilon},\Sigma_{1:k_{x}}^{2,x}V_{1,\epsilon},T_{4,1:k_{y}}V_{1,\epsilon}]\right)$ ;
@@ 28               Exit loop;
3.6 Computational Complexity
We analyze next the computational complexity of Algorithm 4 for backward Euler. For conciseness, we consider the same spatial resolution per dimension, i.e., $N=N_{1}=N_{2}$. Starting from an initial condition $U_{0}S_{0}V_{0}^{\top}$, where $U_{0}\in\mathbb{R}^{N\times r_{0}}$, $V_{0}\in\mathbb{R}^{N\times r_{0}}$, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$, the $m$-th xKrylov iteration generates $U_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, $V_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, and $S_{1}\in\mathbb{R}^{r_{\kappa}\times r_{\kappa}}$, where $r_{\kappa}=(2+\mathcal{R}_{\text{ranks}})r^{\ast}_{0}+r^{(m-1)}$. Here, $r^{(m-1)}$ denotes the basis size from the previous xKrylov iteration (initialized with $r^{(0)}=r_{0}$), and $r^{\ast}_{0}\leq r_{0}$ is the possibly truncated initial rank as produced by line (2) of Algorithm 2. Using SVD-truncated QR, this dimension is then reduced to smaller ranks, $r_{x}$ and $r_{y}$. For simplicity, let $\tilde{r}=\max(r_{x},r_{y})$. The matrix $S_{1}$ from the $m$-th inner iteration solve, initially of size $r_{x}\times r_{y}$, is truncated to rank $r_{1}\leq\min(r_{x},r_{y})$ using Algorithm 3. To simplify the complexity analysis, let $r=\max(r_{0},r^{(m-1)},\tilde{r},(\mathcal{R}_{\text{ranks}}+1)r_{1})$, which we consider independent of the mesh-refinement $N$ for practical purposes. The computational complexity of each step is then estimated as follows:
@@ 3.6 Computational Complexity
Constructing the generalized Krylov basis as outlined in lines (4) and (4) of Algorithm 4 requires performing the operations $A_{1,1:r_{x}}^{-1}U_{0}$ and $A_{3,1:k_{x}}^{-1}U_{0}$, and $A_{2,1:r_{y}}^{-1}V_{0}$ and $A_{4,1:k_{y}}^{-1}V_{0}$, which require solving the systems $AX=U$ and $AX=V$ (subscripts have been omitted for simplicity). With 1D finite-differences, $A$ has a tridiagonal structure, which can be efficiently factorized using direct-solver techniques such as the Thomas algorithm, with $\mathcal{O}(N)$ complexity per column of $X$ [34]. The same complexity holds for applying $P_{1}^{-1}$ and $P_{2}^{-1}$. Additionally, applying $P_{1}U_{0}$, $\Phi_{1,3}U_{0}$, $\Sigma_{1,3}U_{0}$, $P_{2}V_{0}$, $\Phi_{2,4}V_{0}$, and $\Sigma_{2,4}V_{0}$, can be done with $\mathcal{O}(Nr^{\ast}_{0})$ computational complexity for sparse diagonal and tridiagonal matrices, leading to a total complexity of $\mathcal{O}(N(2+\mathcal{R}_{\text{ranks}})r^{\ast}_{0})$ to augment the xKrylov subspaces. The use of Algorithm 2 to compute the orthonormal truncated bases $U_{1}$ and $V_{1}$ incurs a complexity of $\mathcal{O}\left(N\,\max\left[\left(2+\mathcal{R}_{\text{ranks}}\right)\,r_{0}^{\ast}\,r^{(m-1)},\;\left(\left(2+\mathcal{R}_{\text{ranks}}\right)\,r_{0}^{\ast}\right)^{2}\right]\right)\sim\mathcal{O}(Nr^{2})$. Generally, we expect $r$ to be much smaller and independent of $N$ when the solution is sufficiently smooth and exhibits a low-rank structure.
@@
The application of the reduced GSE operator using matrix-matrix multiplication and summation in the Krylov method (GMRES) features a complexity of $\mathcal{O}\left(\tilde{r}^{d+1}\right)$. Applying the ACS preconditioner requires solving an SE of size $r_{x}\times r_{y}$, which can be achieved using the Bartels-Stewart algorithm [15] with a complexity of $\mathcal{O}(\tilde{r}^{d+1})$. Hence, the $S$-step has $\mathcal{O}(r^{d+1})$ complexity.
@@
The QR decomposition to obtain $R_{U}$ and $R_{V}$ in lines (11)-(4) of Algorithm 4 features a complexity of $\mathcal{O}\left(N\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{2}\right)$. The residual computation in line (4) is of $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{d+1}\right)$ due to the matrix multiplications, with an additional $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{d}\right)$ for the Frobenius norm.
It follows that, as advertised, the overall complexity of the algorithm scales as $Nr^{2}$ and as $r^{d+1}$, with $d$ the dimensionality, which will be numerically verified in Section 4.
@@ 4 Numerical Experiments
We demonstrate the overall efficacy of the proposed algorithm with several 2D and 3D numerical examples. We assume homogeneous Dirichlet boundary conditions for simplicity; other boundary conditions (.e.g, periodic, Neumann, etc.) can be easily incorporated in the formulation through the operators $T_{1:4}$ (10). All simulations are performed using MATLAB. For efficient computational handling in 3D, we employ tools from the Tensor Toolbox [3] and the Htucker package [25]. Throughout this section, we denote the diffusion Courant number and the advection Courant-Friedrichs-Lewy (CFL) number by $\lambda_{D}$ and $\lambda_{A}$, respectively, defined as:
$\lambda_{D}=\frac{\Delta t\phi_{\text{max}}}{\Delta x^{2}},\qquad\lambda_{A}=\frac{\Delta t\sigma_{\text{max}}}{\Delta x},$ (34)
where $\phi_{\text{max}}$ and $\sigma_{\text{max}}$ are the maximum of the diffusion and advection coefficients, respectively. These numbers are measures of the condition number of the associated GSEs. For all simulations presented in this section, we did not need to set a maximum number of outer xKrylov iterations because the algorithm consistently converged within a small number of iterations.
@@
and a rank-3 variable diffusion coefficient:
@@
The initial condition is a rank-2 Gaussian:
@@
We use a spatial grid with $N=300$ in all dimensions, and the time-step $\Delta t$ is set to $\lambda_{D}\Delta x^{2}$, with $\lambda_{D}$ ranging from $140$ to $1125$. We test our adaptive-rank algorithm using three different integrators: backward Euler (BE), DIRK2, and DIRK3, as detailed in Tables 3 and 3 in the Appendix. The residual tolerance is $\epsilon_{\texttt{tol}}=10^{-4}$ for BE, DIRK2, and DIRK3, the basis truncation threshold is $\epsilon_{\kappa}=10^{-6}$, the SVD truncation threshold is $\epsilon=10^{-6}$, and the GMRES acceptance threshold is $\epsilon_{\texttt{GMRES}}=10^{-6}$.
We investigate numerically the following: (1) high-order temporal convergence, (2) effectiveness of the xKrylov iterations in capturing the range of solution, (3) performance of the proposed ACS preconditioner, and (4) verification of the overall computational complexity.
High-Order Temporal Convergence: Figure 1(a) illustrates the $L_{1}$ error norm vs time stepping size (characterized by $\lambda_{D}$) by showcasing the performance of the adaptive-rank integrators for BE, DIRK2, and DIRK3. The expected asymptotic temporal error order of convergence for our proposed adaptive-rank integrators, depicted by lines with markers, is observed. Additionally, the algorithm adeptly captures the rank evolution of the solution in the x, y, and z directions, as illustrated in Figure 1(b).
@@
Effectiveness of the xKrylov Iterations: We evaluate the xKrylov subspace dimension required to achieve tolerances of $\epsilon_{\texttt{tol}}=10^{-4},10^{-5},10^{-6}$. We perform a single time-step integration with a fixed step size $\Delta t=10^{-3}$ with grid refinements in $N$. To isolate the effect of $\lambda_{D}$ (a measure of the GSE condition number) on the convergence behavior of xKrylov, we mitigate other potential influencing factors, such as excessive truncation of the Krylov subspace, insufficient convergence of the GMRES-ACS solver, and truncation errors in the constructed solution, by setting related thresholds to $\epsilon=\epsilon_{\texttt{GMRES}}=\epsilon_{\kappa}=10^{-10}$. Figure 2 illustrates the xKrylov subspace dimension required for convergence as a function of the diffusion Courant number $\lambda_{D}$ for a fixed $\Delta t$ and varying spatial resolutions $N$. Our results suggest that, for a fixed tolerance, the convergence of the Krylov subspaces remains largely insensitive to grid refinement across a broad range of $\lambda_{D}$. However, for extremely fine mesh resolutions, beyond typical practical computational needs, the convergence rate of xKrylov scales as $\lambda_{D}^{1/4}$, consistently with the $\kappa^{1/4}$ scaling (with $\kappa$ the condition number) previously reported for the Poisson equation [24, 4].
@@
Performance of ACS-Preconditioner: We compare the performance of the GMRES solver with and without the ACS preconditioner for the inner solve. Figure 3 illustrates the GMRES residual history over iterations for ACS-preconditioned and unpreconditioned cases. Results at the first time step ($\Delta t=0.01$) and the third outer-basis augmentation are shown, comparing different mesh sizes. The results confirm that ACS-preconditioned GMRES delivers mesh-independent and super-linear convergence rates, which is not the case for unpreconditioned GMRES.
@@
Figure 4 (a) illustrates the average condition number of the Galerkin-projected diffusion and advection operators as the dimension of the proposed xKrylov subspaces increases across 12 outer iterations. For these results, the grid size is fixed at $N=1000$ in 3D, and a single time step of $\Delta t=0.01$ is used. Figures 4 (b) - (d) show the GMRES residual histories for the ACS-preconditioned and unpreconditioned cases, corresponding to the second, fourth, and twelfth outer xKrylov iterations. These correspond to xKrylov subspace dimensions of $r=37$, $r=73$, and $r=217$, respectively. As the dimension of the xKrylov subspace grows, unpreconditioned GMRES convergence for the projected problem becomes progressively slower, as expected. In contrast, the ACS-preconditioner effectively maintains rapid convergence rates independent of subspace size $r$.

### v3

@@ 3.4 Truncation and Residual Computation
Following [14], we compute the residual using the identity:
@@
$={\left\|R_{U}\begin{bmatrix}-\tilde{B}_{1}&\mathbf{0}&\mathbf{0}\\ \mathbf{0}&{S_{1,\epsilon}}&\mathbf{0}\\ \mathbf{0}&\mathbf{0}&-\Delta t\,(I_{\mathcal{R}_{\text{ranks}}}\otimes{S_{1,\epsilon}})\end{bmatrix}R_{V}^{\top}\right\|}\leq\epsilon_{\texttt{tol}},$ (23)
@@
[…]
@@
We note that, as discussed in [33] and analyzed in detail later in this study, forming, storing, and evaluating the norm of the residual tensor in (23) can be a very expensive operation when the number of GSE terms, $\mathcal{R}_{\text{ranks}}$, becomes large. This is so because the complexity of a naïve implementation of the residual evaluation scales as $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d+1})$ and the memory storage scales as $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d})$ (although the latter has been reduced in our implementation to $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{2})$ by careful algorithmic choreography). While this is not a pressing issue in our numerical experiments (which feature small-to-moderate values of $\mathcal{R}_{\text{ranks}}$), it could become a limiting factor in applications where $\mathcal{R}_{\text{ranks}}$ becomes large. In such instances, it may be necessary to use alternate strategies as proxy measures of convergence such as the difference between consecutive iterates [33], which is commensurate in cost with the rest of our algorithm.
@@ 3.5 Generalization to Multi-rank Coefficients
As before, we find $\mathscr$$\tilde${L}} by averaging the diagonal terms $(\Phi_{i}^{2,x},\Phi_{j}^{1,y},\Sigma_{k}^{2,x},\Sigma_{l}^{1,y})$ from (25) to arrive at the approximated SE:
@@
We summarize the pseudo-algorithm for multi-rank coefficients in Algorithm 4. The tolerances $\epsilon_{\kappa}$, $\epsilon_{\texttt{GMRES}}$, $\epsilon$, $\epsilon_{\texttt{tol}}$ in the algorithm are set as follows. First, we choose the residual tolerance for xKrylov outer iterations, $\epsilon_{\texttt{tol}}$. To prevent rapid growth of the xKrylov subspaces, we choose the Krylov basis truncation threshold as $\epsilon_{\kappa}\leq\epsilon_{\texttt{tol}}$, typically two to three orders of magnitude smaller. The GMRES-ACS tolerance for the inner reduced system solve, $\epsilon_{\texttt{GMRES}}$, is initially set to be less than $\epsilon_{\texttt{tol}}$ for the first outer-basis iteration. It is subsequently adjusted to be two to three orders of magnitude smaller than the relative residual error from the previous xKrylov outer iteration. This adjustment ensures that errors from the inner solve do not pollute the accuracy of the outer solve. The SVD truncation threshold $\epsilon$ is chosen to be one to three orders of magnitude smaller than the tolerance $\epsilon_{\texttt{tol}}$. Since solution truncation occurs before residual evaluations, the truncation error in the solution should remain below the residual tolerance to prevent residual pollution by over-truncation.
@@ 7    Orthogonalize and truncate to tolerance $\epsilon_{\kappa}$;
17    Compute $\{\_,R_{U}\}=\texttt{QR}\left([U_{1},U_{1,\epsilon},T_{1,1:\ell_{x}}U_{1,\epsilon},\Phi_{1:\ell_{y}}^{1,y}U_{1,\epsilon},T_{3,1:k_{x}}U_{1,\epsilon},\Sigma_{1:k_{y}}^{1,y}U_{1,\epsilon}]\right)$ and
18    $\{\_,R_{V}\}=\texttt{QR}\left([V_{1},V_{1,\epsilon},\Phi_{1:\ell_{x}}^{2,x}V_{1,\epsilon},T_{2,1:\ell_{y}}V_{1,\epsilon},\Sigma_{1:k_{x}}^{2,x}V_{1,\epsilon},T_{4,1:k_{y}}V_{1,\epsilon}]\right)$ ;
@@ 28        Exit loop;
3.6 Analysis of Computational and Memory Complexity
We analyze next the computational complexity and memory storage of Algorithm 4 for backward Euler. For conciseness, we consider the same spatial resolution per dimension, i.e., $N=N_{1}=N_{2}$. Starting from an initial condition $U_{0}S_{0}V_{0}^{\top}$, where $U_{0}\in\mathbb{R}^{N\times r_{0}}$, $V_{0}\in\mathbb{R}^{N\times r_{0}}$, and $S_{0}\in\mathbb{R}^{r_{0}\times r_{0}}$, the $m$-th xKrylov iteration generates $U_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, $V_{1}\in\mathbb{R}^{N\times r_{\kappa}}$, and $S_{1}\in\mathbb{R}^{r_{\kappa}\times r_{\kappa}}$, where $r_{\kappa}=(2+\mathcal{R}_{\text{ranks}})r^{\ast}_{0}+r^{(m-1)}$. Here, $r^{(m-1)}$ denotes the basis size from the previous xKrylov iteration (with $r^{(0)}=r_{0}$), and $r_{0}^{\ast}\leq r_{0}$ is the (possibly truncated) initial rank produced at line (2) of Algorithm 2. While not used in this work, one may impose a maximum allowable rank $r_{\max}$ in the orthogonalization step of Algorithm 1 to prevent the subspace from exceeding available computational resources. Using SVD-truncated QR, this dimension is then reduced to smaller ranks, $r_{x}$ and $r_{y}$. Let $\tilde{r}=\max(r_{x},r_{y})$. The matrix $S_{1}$ from the $m$-th inner iteration solve, initially of size $r_{x}\times r_{y}$, is truncated to rank $r_{1}\leq\min(r_{x},r_{y})$ using Algorithm 3. To simplify the complexity analysis, let $r=\max(r_{0},r^{(m-1)},\tilde{r},(\mathcal{R}_{\text{ranks}}+1)r_{1})$, which we consider independent of the mesh-refinement $N$ for practical purposes. The computational complexity of each step is then estimated as follows:
@@ 3.6 Analysis of Computational and Memory Complexity
Constructing the generalized Krylov basis as outlined in lines (4) and (4) of Algorithm 4 requires performing the operations $A_{1,1:r_{x}}^{-1}U_{0}$ and $A_{3,1:k_{x}}^{-1}U_{0}$, and $A_{2,1:r_{y}}^{-1}V_{0}$ and $A_{4,1:k_{y}}^{-1}V_{0}$, which require solving the systems $AX=U$ and $AX=V$ (subscripts have been omitted for simplicity). With 1D finite-differences, $A$ has a tridiagonal structure, which can be efficiently factorized using direct-solver techniques such as the Thomas algorithm, with $\mathcal{O}(N)$ complexity per column of $X$ [37]. The same complexity holds for applying $P_{1}^{-1}$ and $P_{2}^{-1}$. Additionally, applying $P_{1}U_{0}$, $\Phi_{1,3}U_{0}$, $\Sigma_{1,3}U_{0}$, $P_{2}V_{0}$, $\Phi_{2,4}V_{0}$, and $\Sigma_{2,4}V_{0}$, can be done with $\mathcal{O}(Nr^{\ast}_{0})$ computational complexity for sparse diagonal and tridiagonal matrices, leading to a total complexity of $\mathcal{O}(N(2+\mathcal{R}_{\text{ranks}})r^{\ast}_{0})$ to augment the xKrylov subspaces. The use of Algorithm 2 to compute the orthonormal truncated bases $U_{1}$ and $V_{1}$ incurs a complexity of $\mathcal{O}\left(N\,\max\left(r_{\kappa}\,r^{(m-1)},\;r_{\kappa}^{2}\right)\right)\sim\mathcal{O}(Nr^{2})$. The memory complexity to store the truncated bases is $\mathcal{O}\left(N\,\left(r^{(m-1)}+r_{\kappa}\right)\right)\sim\mathcal{O}(Nr)$. Generally, we expect $r$ to be much smaller and independent of $N$ when the solution is sufficiently smooth and exhibits a low-rank structure.
@@
The application of the reduced GSE operator using matrix-matrix multiplication and summation in the Krylov method (GMRES) features a complexity of $\mathcal{O}\left(\tilde{r}^{d+1}\right)$. Applying the ACS preconditioner requires solving an SE of size $r_{x}\times r_{y}$, which can be achieved using the Bartels-Stewart algorithm [15] with a complexity of $\mathcal{O}(\tilde{r}^{d+1})\sim\mathcal{O}(r^{d+1})$. Hence, the computation of the $S$-step has $\mathcal{O}(r^{d+1})$   complexity. The memory complexity to store the S matrix/tensor is $\mathcal{O}\left(\tilde{r}^{d}\right)\sim\mathcal{O}(r^{d})$.
@@
The QR decomposition to obtain $R_{U}$ and $R_{V}$ in lines (11)–(4) of Algorithm 4 has a complexity of $\mathcal{O}\left(N\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{2}\right)\sim\mathcal{O}(Nr^{2})$. The residual computation in line (4) requires $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{d+1}\right)\sim\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d+1}r^{d+1})$ due to matrix multiplications, with an additional $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{d}\right)\sim\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d}r^{d})$ for the Frobenius norm. The memory storage requirements for the residual tensor may be quite prohibitive if stored fully [it would scale as $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d}r^{d})$]. Instead, we store only the $R_{U}$ and $R_{V}$ factor matrices and the $S$ matrix, and build the residual for the norm calculation on the fly still with computational complexity $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{d+1}r^{d+1})$. Storing $R_{U}$ and $R_{V}$ requires $\mathcal{O}\left(\left(\tilde{r}+(\mathcal{R}_{\text{ranks}}+1)r_{1}\right)^{2}\right)\sim\mathcal{O}(\mathcal{R}_{\text{ranks}}^{2}r^{2})$, and the residual tensor blocks in Eq. 23 require $\mathcal{O}(r^{d})$, resulting in a total storage of $\mathcal{O}(\mathcal{R}_{\text{ranks}}^{2}r^{2})+\mathcal{O}(r^{d})$. Nevertheless, it is clear from this analysis that the residual evaluation may become prohibitively expensive when $\mathcal{R}_{\text{ranks}}\gg 1$. In such situations, the difference between successive solution iterates may be used as a alternate measure of convergence [33], commensurate in cost to the rest of the algorithm.
It follows that the overall complexity of the algorithm scales as $\mathcal{O}(Nr^{2})+\mathcal{O}(r^{d+1})$, and the memory complexity scales to leading order as $\mathcal{O}(Nr)+\mathcal{O}(r^{d})$, with $d$ the dimensionality. Both claims will be numerically verified in Section 4.
@@ 4 Numerical Experiments
We demonstrate the overall efficacy of the algorithm with several 2D and 3D numerical examples. We assume homogeneous Dirichlet boundary conditions for simplicity; other boundary conditions (e.g., periodic, Neumann, etc.) can be easily incorporated in the formulation through the operators $T_{1:4}$ (9). All simulations are performed using MATLAB. For efficient computational handling in 3D, we employ tools from the Tensor Toolbox [3] and the Htucker package [27]. High-order temporal integration using DIRK is used throughout the numerical section. Details on the the DIRK scheme implementation in our adaptive-rank algorithm are provided in Appendix C. Throughout this section, we denote the diffusion Courant number and the advection Courant-Friedrichs-Lewy (CFL) number by $\lambda_{D}$ and $\lambda_{A}$, respectively, defined as:
$\lambda_{D}=\frac{\Delta t\phi_{\text{max}}}{\Delta x^{2}},\hskip 18.49988pt\lambda_{A}=\frac{\Delta t\sigma_{\text{max}}}{\Delta x},$ (34)
where $\phi_{\text{max}}$ and $\sigma_{\text{max}}$ are the maximum of the diffusion and advection coefficients, respectively. The numbers $\lambda_{D}$ and $\lambda_{A}$ are appropriate measures of the maximum eigenvalue of the associated GSE, and by extension of its condition number since the minimum eigenvalue is of $\mathcal{O}(1)$. For all simulations presented in this section, we did not need to set a maximum number of outer xKrylov iterations because the algorithm consistently converged within a small number of iterations.
@@
Note that this flow is divergence-free and therefore a pure vorticity flow, which results in interesting flow dynamics such as differential rotation. For diffusion, we specify a rank-3 variable coefficient as follows:
@@
These diffusion coefficients are relatively balanced and therefore mostly isotropic. We will investigate anisotropic diffusion in future work. The initial condition is a rank-2 Gaussian:
@@
We use a spatial grid with $N=300$ in all dimensions, and the time-step $\Delta t$ is set to $\lambda_{D}\Delta x^{2}$, with $\lambda_{D}$ ranging from $140$ to $1125$. We test the adaptive-rank algorithm using three different integrators: backward Euler (BE), DIRK2, and DIRK3, as detailed in Tables 4 and 4 in Appendix C. The residual tolerance is $\epsilon_{\texttt{tol}}=10^{-4}$ for BE, DIRK2, and DIRK3, the basis truncation threshold is $\epsilon_{\kappa}=10^{-6}$, the SVD truncation threshold is $\epsilon=10^{-6}$, and the GMRES acceptance threshold is $\epsilon_{\texttt{GMRES}}=10^{-6}$.
We investigate numerically the following: (1) high-order temporal convergence, (2) effectiveness of the xKrylov iterations in capturing the range of solution, (3) performance of the ACS preconditioner to solve the reduced GSE, and (4) verification of the overall computational complexity.
High-Order Temporal Convergence: Figure 1(a) illustrates the $L_{1}$ error norm vs time stepping size (characterized by $\lambda_{D}$) by showcasing the performance of the adaptive-rank integrators for BE, DIRK2, and DIRK3. The expected asymptotic order of convergence of the temporal error for the adaptive-rank integrators, depicted by lines with markers, is observed. Additionally, the algorithm adeptly captures the rank evolution of the solution in the $x$, $y$, and $z$ directions, as illustrated in Figure 1(b).
@@
Effectiveness of the xKrylov Iterations: We evaluate the xKrylov subspace dimension required to achieve tolerances of $\epsilon_{\texttt{tol}}=10^{-4},10^{-5},10^{-6}$. We perform a single time-step integration with a fixed step size $\Delta t=10^{-3}$ with grid refinements in $N$. The convergence bound for the extended Krylov method depends on the condition number of the advection-diffusion operator, $\kappa$. Assuming that the diffusion operator dominates the condition number, and that the variable coefficients perturb its eigenvalues only by factors of order unity, $\lambda_{D}$ provides a good estimate of the condition number (i.e., $\kappa=\mathcal{O}(\lambda_{D})$ in the case of the diffusion equation), and is therefore a good figure of merit to assess the convergence of the xKrylov subspaces. As reported in [25], the convergence factor for the extended Krylov subspace method applied to an equation with Kronecker product structure is $\kappa^{1/4}$ in 2D and $\kappa^{2/3}$ in 3D.
To isolate the effect of $\lambda_{D}$ on the convergence of xKrylov, we mitigate other potential influencing factors, such as excessive truncation of the Krylov subspace, insufficient convergence of the GMRES-ACS solver, and truncation errors in the constructed solution, by setting related thresholds to $\epsilon=\epsilon_{\texttt{GMRES}}=\epsilon_{\kappa}=10^{-10}$. Figure 2 illustrates the xKrylov subspace dimension required for convergence as a function of the diffusion Courant number $\lambda_{D}$ for a fixed $\Delta t$ and varying the spatial 1D resolution, $N$. Our results suggest that, for a fixed tolerance, the convergence of the Krylov subspaces remains largely insensitive to grid refinement across a broad range of $\lambda_{D}$. For extremely fine 1D mesh resolutions, $N\sim 10^{4}$, beyond typical practical computational needs, the convergence rate of xKrylov scales as $\lambda_{D}^{2/3}$, consistently with the theoretically expected $\kappa^{2/3}$ scaling in 3D [25, 4] .
@@
Performance of ACS-Preconditioner: Ill-conditioning for the projected system stems from both grid resolution (the projected system becomes more ill-conditioned as we refine the mesh) and rank growth (the projected system becomes larger in size). For effectiveness, the preconditioner must be able to deal with both sources of ill-conditioning. We consider first the scaling of the ACS-preconditioned GMRES inner solve with respect to the 1D resolution $N$. For this, we compare performance with and without the ACS preconditioner at the first time step ($\Delta t=0.01$) and the third outer-basis augmentation (i.e., keeping the subspace size constant, and therefore the rank), varying the mesh size. Figure 3 illustrates the GMRES residual history over iterations for ACS-preconditioned and unpreconditioned cases. The results confirm that ACS-preconditioned GMRES delivers mesh-independent and super-linear convergence rates, which is not the case for unpreconditioned GMRES.
@@
We consider next the performance of the GMRES inner solve with respect to the solution rank $r$. For this test, the grid size is fixed at $N=1000$ in 3D, and a single time step $\Delta t=0.01$ is used. Figure 4 (a) illustrates the average condition number of the Galerkin-projected diffusion and advection operators as the dimension of the xKrylov subspaces increases across 12 outer iterations, confirming that the projected system gets more ill-conditioned as the rank increases. Figures 4 (b) - (d) show the GMRES residual histories for the ACS-preconditioned and unpreconditioned cases, corresponding to the second, fourth, and twelfth outer xKrylov iterations. These correspond to xKrylov subspace dimensions of $r=37$, $r=73$, and $r=217$, respectively. As $r$ grows, unpreconditioned GMRES converges progressively slower, as expected. In contrast, the ACS-preconditioner maintains rapid convergence rates independent of subspace size $r$.
To assess the relative cost of the preconditioner, we measure the ratio of the wall-clock timings for a single GMRES iteration with and without preconditioning for the different solution ranks considered in Fig. 4, yielding $1.64$, $2.09$, and $1.85$, respectively. A preconditioned GMRES iteration is only about twice as expensive as an unpreconditioned one, but results in a dramatic improvement on GMRES’ convergence rate.